OpenStudy (anonymous):

x-intercepts of f(x)=x^2-25 divided by 6+x^4? A medal awaits.....

OpenStudy (mathmale):

x-intercepts are relatively easy. In this case, O-Gal, just set your function f(x) = to zero and solve the resulting equation for x. The great news is that you can completely ignore the denominator in this case. Try solving f(x) = 0 now, OK?

OpenStudy (mathmale):

Solve: x^2-25=0.

OpenStudy (anonymous):

(x+5) (x-5)=0 x+5=0 x-5=0 x=-5, x=5

OpenStudy (mathmale):

Right! Now, would you please write these results as points on the x-axis? Example: (X, Y).

OpenStudy (anonymous):

(-5,0) (5,0)

OpenStudy (anonymous):

thank you

OpenStudy (mathmale):

Very nice! That's it! Notice that when you set out to find vertical asymptotes, the math is essentially the same, except that you're setting the denominator equal to zero and solving for x. Best wishes!